OpenStudy (96mertcank96):

There are five girls and two boys if two people are at randomly chosen what is the probability that the first person is a girl and the second a boy?

OpenStudy (kropot72):

The probabilitgy of choosing a girl first is 5/7. If a girl has been choosen first the probability of choosing a boy next is 2/6. \[P(girl\ then\ boy) = \frac{5}{7}\times \frac{2}{6}=?\] If necessary, simplify the result.

OpenStudy (96mertcank96):

thank you!

OpenStudy (96mertcank96):

i multiply across or diagonally ?

OpenStudy (kropot72):

You're welcome :) \[\frac{5}{7}\times \frac{2}{6}=\frac{5\times 2}{7\times 6}\]

OpenStudy (96mertcank96):

thank you again

OpenStudy (kropot72):

You're welcome :)